Proposal:
K Kids
Target of Oportunity:
No
Time Observed:
0.0
Time Engineering:
0.0
Time lost technical:
11.5
Time lost weather:
0.0
Time lost other:
0.0
Primary Investigator:
Todd Henry
Email:
Institution:
SMARTS
Instrument:
CTIO 1.5m Echelle
Organization:
SMARTS
Total Program Time:
11.5
- Found both the telescope and dome pointing away from the zenith. After recovering those positions found that, anyway, the Linux/TCS machine is down.
- CHIRON afternoon calibrations failed after the file number 54 (out of the 116 expected.) CHIRON was FROZEN, restarted it.
- Restarted CHIRON, did some tests in Slicer as well as in Fiber mode, the system seems to be fine (check files test*fits.) However number of 'counts' (EMAVGSQ) a bit unstable.
- Left the morning calibrations (CALIB2) running as a way to stress / test the CHIRON system.
- DIMM and / or RINGSS are (apparently) down.